OpenStudy (anonymous):

The midpoint of MN is point P at (–4, 6). If point M is at (8, –2), what are the coordinates of point N? A (–16, 14) B (–10, 5) C (2, 2) D (6, 2)

pooja195 (pooja195):

Hi Welcome to openstudy :-) Use the mid-point formula \[\huge~\rm~\frac{ x_2+x_1 }{ 2},\frac{ y_2+y_1 }{ 2 } \]

OpenStudy (anonymous):

THANKS

pooja195 (pooja195):

yw ^_^ let me know what you get :)

OpenStudy (anonymous):

Step 1 : Draw a line and plot the points on the line. Step 2 : Find the point N. Formula for mid-point is\[\huge{(x,y)=\frac{ x_1+x_2 }{ 2 },\frac{ y_1+y_2 }{ 2 }}\]like what @pooja195 said Given us the point M(8,-2) and the mid-point for point P(-4,6)\[Let~say~point~M(8,-2)=(x_1,y_1)\]\[Let~say~point~P(-4,6)=(x,y)\]\[Let~say~point~N(?,?)=(x_2,y_2)\]Now substitue the value into the mid-point formula to find point N\[\huge{(-4,6)=(\frac{ 8+\color{red}{x_2} }{ 2 },\frac{ -2+\color{red}{y_2} }{ 2 })}\]Remember that\[\huge{(\color{red}{x},\color{green}{y})=(\color{red}{\frac{ x_1+x_2 }{ 2 }},\color{green}{\frac{ y_1+y_2 }{ 2 }})}\]\[\huge{\color{red}{x}=\color{red}{\frac{ x_1+x_2 }{ 2 }}}\]\[\huge{\color{green}{y}=\color{green}{\frac{ y_1+y_2 }{ 2 }}}\]Now,back to the question\[\huge{(-4,6)=(\frac{ 8+\color{red}{x_2} }{ 2 },\frac{ -2+\color{red}{y_2} }{ 2 })}\]solve for x_2\[\huge{-4=\frac{ 8+\color{red}{x_2} }{ 2 }}\]solve for y_2\[\huge{6=\frac{ -2+\color{red}{y_2} }{ 2 }}\]

OpenStudy (anonymous):

@giqqles123

OpenStudy (anonymous):

Btw,\[\huge{Welcome~To~OpenStudy!!!}\] @giqqles123